OpenStudy (anonymous):

Find the point of intersection of f(x) and the line y=-9x+6

OpenStudy (anonymous):

you haven't given us f(x)

OpenStudy (anonymous):

oh its the same as the other one..sorry its f(x)=-3x^2-6x+24

OpenStudy (anonymous):

-9x + 6 = -3x^2 - 6x + 24 3x^2 - 3x - 18 = 0 x^2 - x - 6 = 0 (x - 3)(x + 2) = 0 x = 3, -2 y = -9*3 + 6 = -27 + 6 = -21 y = -9*-2 + 6 = 18 + 6 = 24 hence (3, -21) and (-2, 24)